OpenStudy (christina):

how do I find the output if the input is1

jimthompson5910 (jim_thompson5910):

Replace every variable that you're inputting that 1 for with 1 and then evaluate the expression like normal.

jimthompson5910 (jim_thompson5910):

For example, say you have x+10 the input is x = 1, so replace x with 1 to get 1+10 = 11, so the output is 11

OpenStudy (christina):

Do you use the recursive definition to solve it?

jimthompson5910 (jim_thompson5910):

no, just evaluate the expression as if it were a normal arithmetic expression

jimthompson5910 (jim_thompson5910):

are you working with recursive structures?

OpenStudy (christina):

thank you

jimthompson5910 (jim_thompson5910):

do you have the actual problem?

OpenStudy (christina):

yes,make a table of the function values for inputs between 0 and 10 then find a closed form definition that agrees with the table use your closed form definition to find the outputs at 103,104, and 245. g(n)= 0 if n=0 g(n-1)+1 if n>0

jimthompson5910 (jim_thompson5910):

if n = 1, then g(1) = g(0) + 1 = 0+1= 1 if n = 2, then g(2) = g(1) + 1= 1+1 =2 if n = 3, then g(3) = g(2)+1 = 2+1 = 3 see the pattern?

OpenStudy (christina):

see a pattern but still confused lol

jimthompson5910 (jim_thompson5910):

it's simply increasing by 1 (since g(n-1)+1 means, literally, take the previous output and add 1 to it to get the new/next output) So plug in 4, you get one more than 3 which is 4 Plug in 17, you get one more than 16 (17) Plug in x, you get x So plug in 103, you get 103, etc...

OpenStudy (christina):

k thanks going to try it, be back in a few and let you know how it's going thank you